One of the bases of the trapezoid is 5 units and the height of it is 12 units. If the area



of the trapezoid is 138 square units, then what is the length of the other base?

1
Expert's answer
2022-01-10T12:33:46-0500

Let "a=" the first base of the trapezoid, "b=" the second base of the trapezoid,

"h=" the height of the trapezoid.

Then


"S=\dfrac{a+b}{2}\cdot h"

Solve for "b"


"b=\dfrac{2S}{h}-a"

Given "a=5units, h=12units, S=138{units}^2."

Substitute


"b=\dfrac{2(138)}{12}-5=18(units)"

The other base of the trapezoid is 18 units.